OpenStudy (anonymous):

Which expression is equivalent to 15(12 – x)? A. 15 – 12x B. 15 • 12 – x C. 15 + 12 – x D. 15 • 12 – 15x WILL GIVE MEDAL

OpenStudy (chosenmatt):

c I think

OpenStudy (igreen):

Multiply 15 to all terms in the parenthesis..what do you get?

OpenStudy (anonymous):

Use the distribution property: \[a(b+c) = ab + ac\]

OpenStudy (midhun.madhu1987):

a(b-c) = a*b - a*c

OpenStudy (igreen):

@chosenmatt Don't give out answers, plus your answer is incorrect.

OpenStudy (chosenmatt):

wait no I think it's b

OpenStudy (anonymous):

@midhun.madhu1987 Congratulations for Green.. :)

OpenStudy (anonymous):

the answer is not c or b

OpenStudy (anonymous):

@chosenmatt Stop doing that.. :)

OpenStudy (anonymous):

Neither d nor a.. :P

OpenStudy (anonymous):

it is d just like water said

OpenStudy (chosenmatt):

what?

OpenStudy (anonymous):

is it b

OpenStudy (midhun.madhu1987):

Thanks @waterineyes :) :)

OpenStudy (anonymous):

you are welcome.. Keep it up..!! :)